# LoDIP: Low light phase retrieval with deep image prior

### v1

@@ III-B LoDIP: Phase Retrieval
In LoDIP this setup is further generalized by incorporating the deep image prior framework. In the traditional deep image prior setup, for recovering an image $\mathbf{X}\in\mathbb{C}^{n\times n}$ from its measurements $\mathbf{Y}\in\mathbb{C}^{m\times m}$ the optmization problem is formulated as:
@@
where $\mathbf{Y}=\mathcal{A}\left(\mathbf{X}\right)+\eta$, $\mathcal{A}$ is the known forward operator and $\eta$ is the noise. The optimization variable $\mathbf{X}$ is re-parameterized with a new function $g_{\mathbf{W}}\left(\mathbf{z}\right)$, generally a convolutional neural network. Here $\mathbf{W}$ are the learnable parameters of the network and $\mathbf{z}$ is a random input seed which is fixed throughout the optimization process. In LoDIP we choose $z$ to be the recorded diffraction pattern to provide a strong relation between input and desired output during the optimization process.
More in specifically, we extend the deep image prior setup to incorporate the constraints coming from the above setup: we include the known static region, the known sample support as well as the relative illumination doses on sample and static regions. Since the different doses on the sample and the static region lead to very different range of pixel values for $\mathbf{X}$ and $\mathbf{U}$ we add a scaling factor ($k$) to the optimization objective (Section III-B) to bring the expected output of the neural network to the range $[0,1]$. This greatly improves the final reconstruction quality. Finally the optimization problem solved by LoDIP is:
$\min_{\mathbf{W}}\;\ell\left(\mathbf{Y},\left|\mathcal{F}\left(k*g_{\mathbf{W}}\left(\mathbf{z}\right)+\mathbf{U}\right)\right|^{2}\right),$
@@
Following the literature on this topic, we have used a U-Net [46] with skip connections and ReLU activation functions. The implementation of LoDIP is very flexible and can accommodate multiple modifications. First, any suitable network architecture can be used in the proposed setup. Second, LoDIP can be modified to work in the multiple measurements setting of the original in-situ CDI paper.This interesting modification is left for a future work. Third, LoDIP can seamlessly incorporate a rough reconstruction of the sample as initialization so it can be effectively combined with other techniques designed to generate high-quality initializations for iterative phase retrieval methods [44]. Fourth, LoDIP can be used in the presence of a probe function by incorporating it in the forward operator (this was done in Section IV-B to produce the results for experimental data whose generation required a probe). This makes LoDIP easily adjustable for diverse experimental conditions in contrast to existing state-of-the-art methods such as GPS [25] which would require major modifications to accommodate different experimental setups. Fifth, the performance of LoDIP is not sensitive to the relative size or relative location of the sample and the static structure, nor to the specific choice of the static structure. A static structure design step such as in [45, 47] has the potential to improve the performance, but is not necessary. In fact in our experiments we obtain accurate reconstructions independently of the choice of the static structure (specifically, for each google image reconstruction we use a different static region). Finally, unlike Fourier holography [48, 47], the proposed method works with different illuminations of the sample and the static region.
@@ IV Experimental Results and Discussion
TABLE I: Quantitative comparison of LoDIP and HIO-stat performance. For each metric, the reported values represent the mean over all the samples in the dataset
@@ IV-A Data
We perform experiments on three kinds of data. First, we create a simulated data [^1] Data has been made available here using a procedure similar to [49]. Natural images are used to create the sample region and the static region for each sample in the dataset. Examples of the generated images can be seen in Fig. 2. Next, we create a simulation images of biological cell using physically accurate simulations of a gold lacey for the static structure and a biological cell for the sample region [^2] Data is publicly available in [1], these can be found in Fig. 4.
Finally, we demonstrate the applicability of LoDIP on experimental diffraction patterns [^2] Data is publicly available in [1] from live glioblastoma cells measured with a 534 nm HeNe laser. The dataset utilized in this study is sourced from the work of Lo et al. [1]. The static structure is a 100 µ pinhole exposed to the same incident illumination as the sample. Unlike the first two datasets above, this data has been collected using a probe. Further information about the generation of simulated data and optical laser data collection can be found in the original paper [1]. .
Experimental setup and data generation
@@
The illumination in photons per pixel has been calculated by dividing this number of incident photons on the entire sample area by the total number of pixels containing the sample. The lighting conditions have been graded in number of incident photons per pixel (Np/px in Fig. 2). Based on the total number of incident photons, Poisson noise has been applied. Additive white Gaussian noise is commonly used as an approximation for Poisson noise. However, this approximation is only valid at high photon counts. Since in this case we work with low photon counts, we directly sample from the Poisson distribution with mean equal to the number of incident photons on a pixel:
$N_{p}/px=\frac{\text{\# incident photons}}{\text{\# pixels}}$ (IV.1)
$s=\frac{Np}{\sum_{i,j}Y_{i,j}}$ (IV.2)
~$\mathbf{\tilde{}}{Y}=Poisson(\mathbf{Y}\cdot s)\cdot{\frac{1}{s}}$ (IV.3)
@@ IV-B Results
For all experiments we compare the relative performance of the proposed method (LoDIP) with other popular methods which can be used in this setup. Specifically, we compare with Hybrid input-output (HIO) [50]; Deep Image Prior [11] (DIP) with no static region information; HIO-stat which is a modification of HIO that uses a static region; and Generalized proximal smoothing (GPS) [25] which is the state of the art method for phase retrieval. HIO and DIP work with diffraction patterns collected without a static region (see column 1 in Fig. 2. Rest of the methods, namely HIO-stat, GPS and LoDIP use the static region, however LoDIP also exploits neural network priors.
@@
The proposed method’s robustness and generality are evaluated using natural images. Two commonly used natural image datasets for testing phase retrieval algorithms were selected: the Set12 dataset [51, 52, 53] and internet-sourced stock images used in [49]. For each image, diffraction patterns were first generated using the forward model outlined in Section III and Section IV-A. Results were presented for two low-dose scenarios: 800 and 80 photons per pixel. We observed that the Set12 images exhibited higher resolution and finer details compared to the stock images, thus we refer to these as high and low resolution set respectively. The complementary qualities of the datasets enable a more comprehensive evaluation of each method’s denoising capability and achievable resolution. For each dataset, we compared the reconstructions of each method across 12 samples, with each sample utilizing a different image for its static structure and sample region.
Following previous works on reconstruction with noisy measurements [51], we measure the reconstruction accuracy for all experiments in this paper using using Peak Signal-to-Noise Ratio (PSNR). Higher values of PSNR indicate better fidelity; Structural Similarity Index (SSIM) [53] which measures the similarity between two images, evaluating luminance, contrast, and structure for enhanced accuracy in assessing perceptual image quality. Again higher values indicate better reconstruction; R-factor ($R_{real}$) [25] which measures the degree of agreement between observed and predicted data, commonly used in the assessment of image reconstruction quality. A smaller R-factor indicates a closer agreement between the true and reconstructed image.
@@
The first two rows of Table I show the results for PSNR, SSIM and $R_{real}$ averaged over the entire test set of 12 images for the low and high resolutions sets. Since DIP and HIO without reference regions did not provide good results in this case we omit those methods here. From the table we can see that LoDIP attains better results across all metrics for both data sets and photon counts. The better performance of LoDIP is especially clear in the low photon counts case where LoDIP accuracy is approximately double the accuracy of the other methods. Finally, as expected, the error is higher for the high resolution set as more fine detail has to be reconstructed.
@@
Fig. 3 displays a comparison of the results of LoDIP, HIO with static structure and GPS for images in the low resolution set (left) and high resolution set (right). To study the robustness of our method with respect to the experimental setup we used a circular mask for low resolution set and a square one for the high resolution set. From this figure we can clearly see that GPS and LoDIP produce the best results (as already shown by the metrics in the table), but that GPS produces images that are more noisy than LoDIP reconstructions, showing that LoDIP has a strong denoising ability. The figure shows also the choice of the mask for the sample region does not influence the reconstruction quality for LoDIP (and for the other methods).
@@
We also note that iterative methods such as HIO-stat and GPS require parameter tuning by experts in the field and often multiple independent runs of the algorithm are necessary to obtain good reconstruction. On the other hand, LoDIP hyper-parameter tuning is minimal and convergence is usually obtained in one run. These experiments show that LoDIP is able to provide accurate and denoised reconstructions for both photon counts while preserving the resolution obtained by state of the art iterative method like GPS.
@@
Computational microscopy represents a major potential application of the proposed method. Following previous works on PR for CDI [1, 8], we evaluated the performance of LoDIP on realistic and physically accurate simulated data for a prototypical live cell. , The static structure and biological cell were simulated as done in [1]. The results in this section use a simulated 20-nm thick gold lacey pattern as the static structure and a simulated cell consisting of a vesicle containing water and protein aggregates.
Fig. 4 shows the reconstructions from the same HIO, DIP, HIO with static region, LoDIP and GPS at both doses (respectively 800 and 80 photons/pixel). The static region is unknown and estimated using the procedure described in Section III-A. Similar to the case of natural images, GPS and LoDIP produce the best reconstructions in both cases with LoDIP providing a less noisy reconstructed image than GPS especially in the low photon counts case. HIO and DIP without reference regions fail in the reconstruction. The last row of Table I show the results for PSNR, SSIM and $R_{real}$ in the low photons counts case. Again LoDIP outperforms the other methods across all metrics.
FRC is a technique used in microscopy [1, 8] to assess the resolution and overall quality of a reconstruction: it calculates the correlation coefficient between the Fourier transforms of two images, measured within concentric rings of varying spatial frequencies. The resolution is assessed based on the drop in correlation at high frequencies, with a threshold of 0.367 (for more details see [54]). Higher FRC values correspond to better resolution. Fig. 5 displays the Fourier Ring Correlation (FRC) for both photon counts cases for LoDIP and GPS. The resolution of the reconstruction obtained by LoDIP is comparable to the one obtained by GPS. This indicates that the denoising capability of LoDIP results in only a minimal loss of fine-scale details.
@@
In practice, image reconstruction without precise support information presents a common and challenging problem. To conclude this section, we investigate the robustness of the proposed LoDIP method when dealing with only approximately known support for the biological cell example in the low photons count case.
In Fig. 6, the second column shows reconstruction at 80 Np/px when accurate size and shape of the support is known. The last two columns show the reconstruction when the size of the known support estimate is 3 times and 4 times the size of the actual support. From these results we can see that LoDIP reconstructions are robust to inaccurate support specification. Widely used phase retrieval algorithms struggle when accurate support information is not available [43, 26]. This can be partially mitigated by employing heuristic subroutine such as shrinkwrap [26] to refine the support estimate every few iterations. While effective, these end up increasing the total number of hyperparameters needed to be tuned. LoDIP does not require any such subroutine for support estimation.
@@
Finally, we perform an experiment to demonstrate the proposed method on experimentally captured diffraction patterns (one of them is shown in Fig. 7 left column). Unlike the previous experiments, the object is complex-valued and the optical setup includes a probe. The data for this section is the one provided in [1]. We note explicitly that in this dataset the static and sample images have the same high dose, so this is not exactly the setup for which our method is designed. In fact, our method is especially advantageous in the low dose setting, as the previous experiments show, while it obtains comparable results as other methods in the high dose setting. Nonetheless, this setup is very close to the one of interest so we included it in this work. Generating a new experimental dataset with different doses on the static and sample images is left for a future work.
@@
The original in-situ CDI method [1] uses 50 diffraction patterns with a fixed time-invariant static structure required in all the images. HIO-stat and LoDIP use only a single diffraction pattern and perform single-instance phase retrieval. This is an advantage over in situ CDI since as in many experimental setups a truly static structure is difficult to ensure because of the data collection procedure (e.g. in collecting tomographic data, which requires shifting or tilting both the sample and the static structure). In this section, the reconstruction given by in situ CDI is used as a proxy for the ground truth. In fact, a comparison between in situ CDI and LoDIP would be unfair since LoDIP uses 50 times less data (only one diffraction pattern instead of 50). A comparison with GPS is not possible for this data since the current version of GPS does not support Fresnel propagation and adding this would require major modifications which are outside of the scope of this work. In these experiments, LoDIP and HIO-stat have been modified to incorporate the known probe function.
@@
Figure 7: Reconstruction of experimental data (Left) Experimentally captured diffraction pattern. (Right) Reconstruction from different methods.The mean of the top five reconstructions from 20 independent runs is shown in the figure. The static and sample images in this example have the same high dose. LoDIP is especially advantageous in the low dose setting, as the previous experiments show, so as expected, in this high dose setting LoDIP obtains comparable results as in situ CDI and HIO in terms of $R_{F}$. However, LoDIP uses a single diffraction pattern to produce a reconstruction of comparable quality as in-situ CDI which uses 50 samples. Moreover, while HIO-stat reconstruction has similar $R_{F}$ as LoDIP, it contains visible artefacts (indicated by the arrows).
@@
Since there is no ground truth available, to evaluate our results we use R-factor in the Fourier domain. This measures the disagreement between the captured diffraction pattern $\mathbf{Y}$ to the Fourier magnitudes $\lvert\mathcal{F}\left(\hat{\mathbf{X}}\right)\rvert$ of the reconstruction $\hat{\mathbf{X}}$ and is defined as:
$R_{F}(\hat{\mathbf{X}})=\frac{\sum_{i,j}\left|\lvert\mathcal{F}\left(\hat{\mathbf{X}}\right)\rvert_{i,j}-\mathbf{Y}_{i,j}\right|}{\sum_{i,j}Y_{i,j}}$ (IV.4)
@@
The reconstruction obtained through in situ CDI yields an R-factor of $30.22\%(\pm 0.97\%)$. As explained before in situ CDI utilizes 50 diffraction patterns for its reconstruction. Comparing this R-factor directly with LoDIP, which only uses a single diffraction pattern, would be unfair due to the substantial difference in data quantity needed to obtain the reconstruction. However, we can use in situ CDI R-factor as a benchmark to assess the performance of our method. The average $R_{F}$ from 20 independent reconstructions based on a single diffraction pattern using both LoDIP and HIO-stat is presented in Table II. Remarkably, LoDIP achieves a comparable R-factor to in-situ CDI, highlighting its efficacy even without the need for multiple diffraction patterns, making it less data-intensive. Interestingly, HIO-stat also demonstrates a comparable R-factor to LoDIP in this scenario. It’s important to note, however, that this success is primarily attributed to the high photon count setting of this experiment, where many methods perform well. It is crucial also to note that, as demonstrated in previous experiments, HIO-stat’s performance diminishes in low photon count settings, providing less accurate results than LoDIP in this more realistic case. Finally, Fig. 7 shows the averaged top 5 reconstructions out of 20 independent runs for all methods. We see that HIO-stat introduces visible artifacts in its reconstruction while LoDIP does not (again the in situ CDI reconstruction is treated as a proxy for the ground truth in this case). Therefore, these results highlight the robustness and adaptability of LoDIP, showcasing its ability to easily accommodate changes in the data acquisition setup and maintain reliable reconstructions across varying experimental conditions.

### v2

@@ III-B LoDIP: Phase Retrieval
In LoDIP this setup is further generalized by incorporating an implicit generative prior. For a given inverse problem aimed at recovering an image $\mathbf{X}\in\mathbb{C}^{n\times n}$ from its measurements $\mathbf{Y}\in\mathbb{C}^{m\times m}$ the implicit generative prior optmization problem is formulated as:
@@
where $\mathbf{Y}=\mathcal{A}\left(\mathbf{X}\right)+\eta$, $\mathcal{A}$ is the known forward operator and $\eta$ is the noise. The optimization variable $\mathbf{X}$ is re-parameterized with a new function $g_{\mathbf{W}}\left(\mathbf{z}\right)$, generally a CNN. Here $\mathbf{W}$ are the learnable parameters of the network and $\mathbf{z}$ is a random input seed which is fixed throughout the optimization process. $g_{\mathbf{W}}$ essentially works as a generative network which is optimized to output a single image.
Further, we extend this formulation to incorporate the constraints coming from the above setup: we include the known static region, the known sample support as well as the relative illumination doses on sample and static regions. Finally the optimization problem solved by LoDIP is:
$\min_{\mathbf{W}}\;\ell\left(\mathbf{Y},\left|\mathcal{F}\left(g_{\mathbf{W}}\left(\mathbf{z}\right)+\mathbf{U}\right)\right|^{2}\right),$
@@
Extensions to LoDIP framework The LoDIP framework is highly versatile and can accommodate a variety of modifications. Firstly, it can incorporate any suitable implicit generative prior based architecture, including variants of DIP [58, 59]. Secondly, LoDIP can be easily tailored to different experimental conditions by integrating them into the forward operator. This adaptability is demonstrated in Section IV-B, where experimental data captured using a probe was used. In this instance, LoDIP was adjusted to include the known probe function, in contrast to existing state-of-the-art methods like GPS [6], which would necessitate significant alterations for different experimental setups.
Moreover, LoDIP can smoothly integrate a rough reconstruction of the sample as initialization, enabling it to be effectively combined with other techniques aimed at generating high-quality initializations for iterative phase retrieval methods [55]. Finally, LoDIP can be conveniently modified to function in the multiple measurements setting of the original in-situ CDI paper.
@@ IV Experimental Results and Discussion
TABLE I: Quantitative comparison of LoDIP and HIO-stat. For each metric, the reported values represent the mean over all the samples in the dataset. The standard deviation reported for PSNR shows the robustness of the method to different random inputs.
@@ IV-A Data
We perform experiments on three kinds of data. First, we create a simulated data [^3] Data is available at https://github.com/raunakmanekar/LoDIP-PR. using a procedure similar to [60]. Natural images are used to create the sample region and the static region for each sample in the dataset. Examples of the generated images can be seen in Fig. 1. Next, using the data provided in [9] we create a simulated image of biological cell using physically accurate simulations of a gold lacey for the static structure and a biological cell for the sample region,these can be found in Fig. 4, 6.
Finally, we demonstrate the applicability of LoDIP on experimental diffraction patterns from live glioblastoma cells (see Fig. 7). In this setup, the collimated beam with a wavelength of 534 $nm$ incident onto a pair of 100 $\mu m$ pinholes with an edge-to-edge separation distance of 100 $\mu m$. This produced a pair of beams incident onto the sample 400 $\mu m$ downstream of the dual pinhole aperture with the same intensity, as no dose reduction was performed in this experiment. The diffraction intensity from the sample exit wave was recorded using a CCD detector placed at the focal point of a lens adjacent to the sample. The static structure is a 100 µ pinhole exposed to the same incident illumination as the sample. Unlike the first two datasets above, this data has been collected using a probe. This dataset is sourced from Lo et al. [9]. Further information about the optical laser data collection can be found in the original paper.
Simulated data generation
@@ IV-B Results
Methods For all experiments we compare the relative performance of the proposed method (LoDIP) with other popular methods which can be used in this setup. Specifically, we compare with Hybrid input-output (HIO) [61]; Deep Image Prior [14] (DIP) with no static region information; HIO-stat which is a modification of HIO that uses a static region; and Generalized proximal smoothing (GPS) [6] which is the state of the art method for phase retrieval. HIO and DIP work with diffraction patterns collected without a static region (see column 1 in Fig. 1. Rest of the methods, namely HIO-stat, GPS and LoDIP use the static region, however LoDIP also exploits neural network priors.
@@
Metrics. Following previous works on reconstruction with noisy measurements [62], we measure the reconstruction accuracy for all experiments in this paper using using Peak Signal-to-Noise Ratio (PSNR). Higher values of PSNR indicate better fidelity; Structural Similarity Index (SSIM) [63] which measures the similarity between two images, evaluating luminance, contrast, and structure for enhanced accuracy in assessing perceptual image quality (higher the better); R-factor ($R_{real}$) [6] which measures the degree of agreement between observed and predicted data, commonly used in the assessment of image reconstruction quality (smaller the better).
@@
Two commonly used natural image datasets for testing phase retrieval algorithms were selected: the Set12 dataset [62, 64, 63] and internet-sourced stock images used in [60]. For each image, diffraction patterns were first generated using the forward model outlined in Fig. 3 and Section IV-A. Results were presented for two low-dose scenarios: 800 and 80 photons per pixel. Different doses correspond in practice to different amounts of Poisson noise. Good reconstruction across dose amounts shows the robustness of the method with respect to different noise levels.
We observed that the Set12 images showed higher resolution and finer details compared to the stock images, hence we distinguish them as the high and low resolution sets, respectively. The complementary characteristics of these datasets allow for a thorough evaluation of each method’s denoising capability and achievable resolution. Across each dataset, we compared reconstructions from all methods using 12 samples, each utilizing a different image for its static structure and sample region.
@@
Table I shows results averaged over the entire test set of 12 images for the low and high resolutions sets. Since DIP and HIO without reference regions did not provide good results in this case we omit those methods here. LoDIP attains better results across all metrics for both data sets and photon counts. The better performance of LoDIP is especially clear in the low photon counts case where LoDIP accuracy is approximately double the accuracy of the other methods. Finally, as expected, the error is higher for the high resolution set as more fine detail has to be reconstructed.
@@
Discussion Fig. 3 displays a comparison of the results of LoDIP, HIO-stat and GPS for images in the low resolution set (left) and high resolution set (right). To study the robustness of our method with respect to the experimental setup we used a circular mask for low resolution set and a square one for the high resolution set. GPS and LoDIP yield the best results, however, GPS generates noisier images compared to LoDIP, highlighting LoDIP’s strong denoising ability. The figure shows also the choice of the mask for the sample region does not influence the reconstruction quality for LoDIP (and for the other methods).
@@
We also note that iterative methods such as HIO-stat and GPS require parameter tuning by experts in the field and often multiple independent runs of the algorithm are necessary to obtain good reconstruction. On the other hand, LoDIP hyper-parameter tuning is minimal and convergence is usually obtained in one run.
@@
Computational microscopy represents a major potential application of the proposed method. Following previous works on PR for CDI [9, 10], we evaluated the performance of LoDIP on realistic and physically accurate simulated data for a prototypical live cell. The static structure and biological cell were simulated as done in [9]. The results in this section use a simulated 20-nm thick gold lacey pattern as the static structure and a simulated cell consisting of a vesicle containing water and protein aggregates. The static region is unknown and estimated using the procedure described in Section III-A.
Discussion Similar to the case of natural images, Fig. 4 shows that GPS and LoDIP produce the best reconstructions in both cases with LoDIP providing a less noisy reconstructed image than GPS especially at 80 Np/px. HIO and DIP without reference regions fail in the reconstruction. The last section of Table I shows that LoDIP outperforms the other methods across all metrics.
Comparing Resolution. Fig. 5 displays the Fourier Ring Correlation (FRC) for both photon counts cases for LoDIP and GPS. Higher FRC values correspond to better resolution. The resolution of the reconstruction obtained by LoDIP is comparable to the one obtained by GPS. This indicates that the denoising capability of LoDIP results in only a minimal loss of fine-scale details.
@@
Reconstruction without precise support PR without precise support information presents a common and challenging problem. From Fig. 6, we can see that LoDIP reconstructions are robust to inaccurate support specification. Widely used phase retrieval algorithms struggle when accurate support information is not available [53, 20] and require additional measures for support estimation[65].
@@
Finally, we demonstrate the proposed method on experimentally captured diffraction patterns (one of them is shown in Fig. 7 left column). Unlike the previous experiments, the object is complex-valued and the optical setup includes a probe. We note explicitly that in this dataset the static and sample images have the same high dose, so this is not exactly the setup for which our method is designed. In fact, our method is especially advantageous in the low dose setting, as the previous experiments show. Nonetheless, this setup is very close to the one of interest so we included it in this work.
@@
Comparison with in-situ CDI. The original in-situ CDI method [9] uses 50 diffraction patterns with a fixed time-invariant static structure required in all the images. HIO-stat and LoDIP use only a single diffraction pattern and perform single-image phase retrieval. This is an advantage over in-situ CDI since as in many experimental setups a truly static structure is difficult to ensure because of the data collection procedure (e.g. in collecting tomographic data, which requires shifting or tilting both the sample and the static structure). In this section, the reconstruction given by in-situ CDI is used as a proxy for the ground truth. In fact, a comparison between in-situ CDI and LoDIP would be unfair since LoDIP uses 50 times less data (only one diffraction pattern instead of 50).
Comparison with GPS. A comparison with GPS is not possible for this data since the current version of GPS does not support Fresnel propagation and adding this would require major modifications which are outside of the scope of this work. In these experiments, LoDIP and HIO-stat have been modified to incorporate the known probe function.
@@
Figure 7: Reconstruction of experimental data (Left) An experimentally captured diffraction pattern. (Right) Reconstruction of two samples from different methods. The mean of the top five reconstructions from 20 independent runs is shown in the figure. The static and sample images in this example have the same high dose. LoDIP is especially advantageous in the low dose setting, as the previous experiments show, so as expected, in this high dose setting LoDIP obtains comparable results as in-situ CDI and HIO in terms of $R_{F}$. However, LoDIP uses a single diffraction pattern to produce a reconstruction of comparable quality as in-situ CDI which uses 50 samples. Moreover, while HIO-stat reconstruction has similar $R_{F}$ as LoDIP, it contains visible artefacts (indicated by the arrows).
@@
Metrics. Since there is no ground truth available, to evaluate our results we use R-factor in the Fourier domain. This measures the disagreement between the captured diffraction pattern $\mathbf{Y}$ to the Fourier magnitudes $\lvert\mathcal{F}\left(\hat{\mathbf{X}}\right)\rvert$ of the reconstruction $\hat{\mathbf{X}}$ and is defined as:
$R_{F}(\hat{\mathbf{X}})=\frac{\sum_{i,j}\left|\lvert\mathcal{F}\left(\hat{\mathbf{X}}\right)\rvert_{i,j}-\mathbf{Y}_{i,j}\right|}{\sum_{i,j}Y_{i,j}}$ (IV.1)
The reconstruction obtained through in-situ CDI yields an R-factor of $30.22\%(\pm 0.97\%)$. As explained before in-situ CDI utilizes 50 diffraction patterns for its reconstruction. Comparing this R-factor directly with LoDIP, which only uses a single diffraction pattern, would be unfair due to the substantial difference in data quantity needed to obtain the reconstruction. However, we can use in-situ CDI R-factor as a benchmark to assess the performance of our method.
@@
Discussion. The average $R_{F}$ from 20 independent reconstructions based on a single diffraction pattern using both LoDIP and HIO-stat is presented in Table II. Remarkably, LoDIP achieves a comparable R-factor to in-situ CDI, highlighting its efficacy even without the need for multiple diffraction patterns, making it less data-intensive. Interestingly, HIO-stat also demonstrates a comparable R-factor to LoDIP in this scenario. It’s important to note, however, that this success is primarily attributed to the high photon count setting of this experiment, where many methods perform well. It is crucial also to note that, as demonstrated in previous experiments, HIO-stat’s performance diminishes in low photon count settings, providing less accurate results than LoDIP in this more realistic case. Finally, Fig. 7 shows the averaged top 5 reconstructions out of 20 independent runs for all methods. We see that HIO-stat introduces visible artifacts in its reconstruction while LoDIP does not (again the in-situ CDI reconstruction is treated as a proxy for the ground truth in this case). Therefore, these results highlight the robustness and adaptability of LoDIP, showcasing its ability to easily accommodate changes in the data acquisition setup and maintain reliable reconstructions across varying experimental conditions.